OpenStudy (deadshot):

What is the simplified form of the quantity 4 x squared minus 25 over the quantity 2 x plus 5 ?

OpenStudy (mertsj):

\[\frac{4x^2-25}{2x+5}\]

OpenStudy (mertsj):

Factor the numerator. Use this: \[a^2-b^2=(a-b)(a+b)\]

OpenStudy (deadshot):

So it's \[(2x-5)(2x+5)?\]

OpenStudy (deadshot):

and then the (2x+5) cancel out, and all that's left is (2x-5), making (2x-5) the answer. Right?

OpenStudy (mertsj):

yes

OpenStudy (deadshot):

Is it (2x - 5, with the restriction x ≠ -five over 2) or ( 2x - 5, with the restriction x ≠ five over 2)?

OpenStudy (mertsj):

x cannot equal -5/2

OpenStudy (deadshot):

Okay, Thanks!

OpenStudy (mertsj):

yw